A researcher is interested in testing two different dosages of a new sleeping medication in a phase II clinical trial against a control group receiving a placebo. Suppose the following table observes the hours of sleep reported the preceding evening for subjects randomly assigned to each of the three groups. α = .05

Dosage One 8.5, 7.9, 8.6, 8.4, 7.6, 9.1.

Dosage Two- 8.1, 7.9 , 7.9, 7.6, 7.4.

Control Group- 5.2, 5.6, 5.7, 5.9, 6.2, 6.4, 6.1.

A) Calculate the MSB and MSW for a one-way analysis of variance procedure.

B) Using the information above, calculate an F statistic, and provide an interpretation of your p-value. (Continue as though all assumptions for ANOVA are met). How would you solve this without using excel?
Answer #1
treatment 1 2 3
count, ni = 6 5 7
mean , x̅ i = 8.350 7.78 5.87
std. dev., si = 0.532 0.277 0.407
sample variances, si^2 = 0.283 0.077 0.166
total sum 50.1 38.9 41.1 130.1 (grand sum)
grand mean , x̅̅ = Σni*x̅i/Σni =   7.23
square of deviation of sample mean from grand mean,( x̅ - x̅̅)² 1.259 0.305 1.840
TOTAL
SS(between)= SSB = Σn( x̅ - x̅̅)² = 7.556 1.525 12.878 21.95883
SS(within ) = SSW = Σ(n-1)s² = 1.415 0.308 0.994 2.717

no. of treatment , k =   3
df between = k-1 =    2
N = Σn =   18
df within = N-k =   15
  
mean square between groups , MSB = SSB/k-1 = 21.9588/2 = 10.9794
  
mean square within groups , MSW = SSW/N-k = 2.717/15 = 0.1812

b)

F-stat = MSB/MSW = 10.9794/ 0.1812 = 60.6087

P value =   0.0000

anova table
SS df MS F p-value
Between: 21.959 2 10.979 60.61 0.0000
Within: 2.717 15 0.181
Total: 24.676 17